OpenStudy (anonymous):

PLEASE HELP ME I WILL MEDAL AND FAN!! Jess observed that 60% of the people at a mall on a particular day shopped for clothes. If 2500 people at the mall did not shop for clothes that day, the number of people who shopped for clothes that day was ______.

OpenStudy (kaisertheslayer):

3750

OpenStudy (anonymous):

Can you explain??

OpenStudy (kaisertheslayer):

yes wait...

OpenStudy (anonymous):

\[\frac{2500}{0.4}=\frac{x}{0.6} \] Solve the above for x.

OpenStudy (kaisertheslayer):

2500/x = 40/100 solve for x, which is the total amount of people once we find x, subtract 2500 from it, and we can find the number of people that shopped

OpenStudy (anonymous):

who was right?? @robtobey i am confused

OpenStudy (anonymous):

@pooja195 can you help?

OpenStudy (anonymous):

anyone??

OpenStudy (anonymous):

yep, 3750

OpenStudy (anonymous):

ok.. thanks :)

OpenStudy (anonymous):

x(0.4)=2500 x=6250 (total amount of people)

OpenStudy (anonymous):

6250 x 0.6 = 3750 (people who shopped for clothes)

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

great